An object experiences three forces (and only three forces): 26N in the positive x-direction, 10N in the negative y-direction, and 72N at an angle of 35∘ above the negative x-axis. What is the magnitude of the net force acting on this object?
Answer #1

here,

F1 = 26 N i

F2 = - 10 N j

F3 = 72 N * ( - cos(35) i + sin(35) j)

F3 = ( - 58.97 i N + 41.3 j N)

the net force , F = F1 + F2 + F3

F = - 32.97 N i + 31.3 j N

the magnitude of the net force acting on this object , |F| = sqrt(32.97^2 + 31.3^2) = 45.5 N